# Milliarcsecond astrometric oscillations in AGN as a precursor of multi-messaging gravitational wave events

### v1

@@ 2 Astrometric oscillations of the quasar J02041514
The source has been observed intensively by the geodetic IVS network for about 15 years until 2009. An apparently anomalous behaviour of J02041514 was reported in the ICRF2 paper (Fey et al., 2015), where this source was referred to as specially-handled target, i.e., requiring a special procedure for obtaining astrometric solutions. This explains its exclusion from the regular IVS observational program after 2009. The available body of VLBI observations of this source enabled us to suspect a non-random behaviour of its astrometric parameters.
Bottom plots: The temporal variations of right ascension (left) and declination (right) of VLBI astrometric measurements centred to the J2000.0 position RA = and DEC = (Xu et al., 2019). The trend with coefficients from Table 1 has been removed. The observational data points are shown in black with their formal error bars. An apparent decrease of data time density after 2009 is due to the recognised anomalous astrometric behaviour and exclusion of the source from regular IVS observing runs (see Section 2 for explanation and reference). The blue curves represent the model fit with parameters listed in Table 1. The shaded blue area around the blue curves indicates the 99.73% confidence interval () for the model prediction.
@@ 2.1 Brief review of astrophysical properties of the source J02041514
The object J02041514 (PKS 0202149, NRAO 91, 4C15.05) is a well-studied flat-spectrum variable radio source associated with AGN – see a concise review of its properties by Jones et al. (2018). Its classification is uncertain. In some works, it is referred to as a blazar (Piner & Kingham, 1998), while others not too reliably classify it as a Seyfert galaxy (Véron-Cetty & Véron, 2006). Its redshift has been recently estimated as (Jones et al., 2018). In the context of the current study, it is important to note that in many VLBI observations at cm wavelengths, the source demonstrates a core–jet morphology with apparent relative superluminal motion of its components with the speed up to (Lister et al., 2013), where denotes the speed of light. A hint on possible transverse motion of a jet component relative to the jet axis is reported on the basis of several VLBI observations by Piner & Kingham (1998). The highest resolution image of the source at 5 GHz obtained with the space VLBI VSOP mission confirms the jet at sub-milliarcsecond scales (Dodson et al., 2008). The source has an appreciable flux density at mm wavelengths (e.g., Owen et al., 1981; Chen & Wright, 2009; Agudo et al., 2014). This, in combination with the astrometric oscillations discussed in the following subsection, makes the object J02041514 a suitable target for future multi-messenger studies discussed in Section 5.
@@ 2.2.1 Generalised Lomb-Scargle periodogram
(1) – the model terms;
(2) and (5) – the estimated coefficients for RA and DEC, respectively;
(3) and (6) – physical dimensions of corresponding coefficients;
(4) and (7) – the corresponding false alarm probability values. The -value indicates the probability of obtaining such the parameter from the observational data under the null hypothesis, i.e., assuming the corresponding term has no effect.
@@
| Term | Estimate | Phys. dimension | -value | Estimate | Phys dimension | -value |
@@
To quantitatively analyse the time series of VLBI astrometric measurements of the source J02041415, we first computed a generalised Lomb–Scargle (GLS) periodogram. The GLS method is widely employed for detecting periodicity in unevenly spaced data (Lomb, 1976; Scargle, 1982); see also VanderPlas (2018) for a comprehensive discussion of its practical applications in the analysis of experimental data. The technique involves fitting a periodic function, , using the (weighted) least-squares method and determining the optimal frequency from a predefined frequency grid that minimises the residuals. In our analysis, we adopted the periodogram formalism developed by Baluev (2014).
Let us consider a time series of measurements taken at times , each associated with an RMS uncertainty . Each observation is thus represented as a triplet (, , ), where the periodicity (if any) is unknown. We consider two competing models for the data: the null (trend-only) model, denoted by , and the alternative model, given by
@@
and is calculated for each trial frequency by fitting the model to the data using the weighted least-squares method.
Several software packages are publicly available for computing the Lomb–Scargle periodogram; however, most of them fit either a purely periodic function or a periodic function with a constant term to the data, without accounting for linear or higher-order trends. As a result, it is often necessary to remove the trend from the data prior to computing the periodogram. Although this approach is not inherently incorrect, it may introduce a slight shift in the estimated period. Therefore, it is generally preferable to fit the periodic component simultaneously with the trend. We computed the periodogram using the Lomb–Scargle algorithm as implemented in the Astropy library (Astropy Collaboration et al., 2013, 2018). Specifically, we used the astropy.timeseries.LombScargle class.
In right ascension, the quadratic trend was found to be significant, whereas the cubic term was not. In contrast, for declination, the cubic trend was significant. Therefore, we removed the corresponding trends before computing the periodogram: the quadratic trend
@@
before computing the periodogram. By default, the fitting model in astropy.timeseries is defined as:
@@ 2.2.2 Statistical significance of the astrometric oscillations
The statistical significance of the highest peak in the PSD can be expressed in terms of the false alarm probability (FAP), which quantifies the likelihood of mistaking random noise for a genuine signal.
One of the most reliable methods for estimating FAP employs the bootstrap approach (Efron, 1979). It evaluates the robustness of model fitting by repeatedly resampling the observed data with replacement and refitting the model to each resample. This procedure enables the assessment of parameter uncertainty (FAP in our case) without relying on strong distributional assumptions. As shown by VanderPlas (2018), this approach might be more efficient than the Monte Carlo simulation in the context of PSD evaluation in the absence of reliable assumptions on the noise distribution. We generated a large number of synthetic datasets by randomly resampling the observed values with replacement, while keeping the observation times fixed. For each resampled dataset, a Lomb–Scargle periodogram was computed, and the maximum peak recorded. The FAP was then estimated as the fraction of bootstrap iterations in which the maximum peak exceeds the peak observed in the original data.
Another approach involves the analytical approximation for the extreme value of FAP, developed by Baluev (2008) using the theory of stochastic processes. This method is computationally efficient and particularly advantageous when analysing large numbers of time series, such as those encountered in large-scale astronomical surveys. It enables fast and reliable estimation of the FAP associated with periodogram peaks, especially when spectral leakage is minimal or can be neglected. Even in the presence of significant aliasing, the approximation remains conservative and does not increase the likelihood of false detections.
@@
| Declination: | FAP, | FAP |
These results suggest that the detected periodicity is statistically significant in declination, but less so in right ascension.
Figure 1 (upper panels, left and right) shows the GLS periodogram of VLBI astrometry right ascension and declination of J02041415, respectively. The highest significant peak in both right ascension and declination occurs at the same period of years. The lower panels of Fig. 1 show the periodic fit in the time series of VLBI astrometry of J02041415. Table 1 summarises the best-fitted amplitudes of trend components and harmonics together with their statistical confidence parameters, obtained in the analysis described above in subsection 2.2.1.
@@ 2.2.3 Validation of the periodicity by cross-spectrum
In order to validate the astrometric oscillation of the source J02041415 described above, we applied to the observational data the cross-spectrum method. It is used to identify periodic signals common to two time series by computing the Fourier transform of their cross-covariance function. The modulus of the cross-spectrum indicates the strength of shared periodic components, in our case – right ascension and declination time series. This technique is particularly effective for detecting coherent periodicities that may be obscured in individual power spectra due to noise or aliasing (Priestley, 1981).
The autocorrelation function is related to the spectral density function through the Fourier transform. Specifically, the spectral density function is given by:
@@
where is the cross-correlation function of two time series, and is the corresponding cross-spectral power density, is the lag.
The cross-spectral density function is generally a complex-valued function and can be represented in two equivalent forms:
where:
and are the real and imaginary parts (or symmetric and antisymmetric components) of the complex function ,
is the amplitude spectrum,
@@
Figure 2 shows the cross-spectrum, revealing a common period in right ascension and declination of years. Although this value differs slightly from the result obtained with the GLS method in subsection 2.2.1, the difference appears to be small and could be a subject of a future detailed investigation.
@@ 3 Qualitative evaluation of the astrometric pattern of J02041514
In the case of sources with complex mas-scale structure, absolute astrometric positions could be affected by changes in the brightness distribution with time. In extreme cases, the effect of relative fading and brightening of components could cause an apparent ‘jump’ up to mas in the radio source position (Titov et al., 2022), (Osetrova et al., 2024). However, generating a persistent harmonic oscillating celestial trajectory this way is very unlikely.
@@
The oscillating pattern of absolute celestial coordinate estimates for the source J02041514 might have a variety of explanations. We must note that VLBI astrometry deals with centroids of the source’s brightness distribution, which are usually associated with the base of jet in AGNs. These are optically thick regions of jets with typically flatter or inverted spectra at radio frequencies. The jet base areas do not coincide with the 3D position of central SMBHs. Thus, what is seen in Fig. 1 is not a motion of SMBHBs themselves. But their motion manifests itself in the apparent motion of the astrometrically detected structural component of the source. A detailed physical description of the causal relationship between the binary system’s orbital period and the observed period of astrometric oscillations associated with a jet-related morphological component warrants a dedicated investigation, which lies beyond the scope of the present paper. However, recent general relativistic magnetohydrodynamic (GRMHD) simulations of binary black hole systems suggest that, depending on the system’s eccentricity, the periodicity of accretion and jet emission ranges from approximately 1.0 to 0.7 times the binary orbital period (Manikantan et al., 2025). Therefore, in the subsequent sections, our assumption that the period of the SMBHB in J02041514 is equal to the observed period of astrometric oscillations introduces no significant error for the toy model considered.
@@
Various effects might be invoked for explanations of the apparent oscillating behaviour of the jet base. One of the most studied cases, involving both long-term flux density variability monitoring in optics and radio, as well as structural studies at VLBI (milliarcsecond) scales, the blazar OJ287, considered to be a likely SMBHB. However, in this and several other similar cases, alternative physical causes of temporal, structural or astrometric manifestations, e.g., the impact of Lense–Thirring effect on an accretion disc or jet precession, are considered as responsible for the observed phenomenology (for a review of the issue see Britzen et al. (2023) and references therein). Among those causes, periodic quasi-orbital motion of the jet base associated with an orbital motion of SMBH in a binary system is at least no less plausible than others, especially coupled with the modality of SMBHB existence in the Universe (see Section 1). Moreover, SMBHB must evolve, and apparent orbital motion of AGN, detected via oscillating behaviour of associated with it bright morphological features in VLBI images of AGN, could be looked at as an indicator of a potential presence of SMBHB.
@@
To certain extent, our interpretation of apparent oscillating astrometric pattern detected in the radio domain is similar to varstrometric detections of potential sub-kpc dual AGNs based on Gaia optical astrometry (Hwang et al., 2020). Another search of potential SMBHBs via their quasi-periodic patterns in optical light curves was reported recently by Chen & Jiang (2024). We also note the recent work by Makarov et al. (2024) in which a novel periodogram analysis algorithm applied to VLBI astrometric monitoring of 259 ICRF3 sources resulted in detection of 49 objects with quasi-periodic celestial position variations, including the sources of our interest, J02041514. They report the detection of a quasi-periodic signal in declination only, with a period of 8.05 years. This value slightly differs from our period estimate, 7.9 years. The reason for this discrepancy requires further investigation. We note, however, that, qualitatively, the absence of a detected quasi-periodic signal in right ascension for J02041514 by Makarov et al. (2024) is consistent with the clearly lower confidence in periodicity detection in right ascension compared to declination in our analysis (see Fig. 1) and FAP values in Table 1 and subsection 2.2.2).
@@
Additional justification of our logic comes from the fact that if an SMBHB evolves all the way toward full coalescence, accompanied by a GW burst, it must end up as a single SMBH. Indeed, some known objects, like the quasar 3C 186 are suspected to be a post-coalescence product of a former SMBHB (Boschini et al., 2024). Another suspected case of merging SMBHs or even a post-coalescence recoiling SMBH is reported recently in the radio quasar PKS 090357 based on optical spectroscopy (Goldoni et al., 2024).
@@
In the following section, we analyse an evolution of SMBHB with astrometric oscillations. Such the binary system will eventually become a mutimessenger with very distinct electromagnetic and GW imprints. We do so for the source discussed in this work, J02041514. This source offers is a straightforward example allowing us to focus on properties of SMBHB as a GW generator. However, the analysis presented in the following section is applicable to other sources with apparent astrometric oscillations, and wider SMBHBs identified by other techniques.
@@ 4 Toy model of SMBHB evolution
As was mentioned in the Introduction, the formation and evolution of SMBHBs encompasses three phases: the pairing phase, the hardening stage, and the final coalescence phase (see Begelman et al., 1980; Colpi, 2014, and references therein). In this section, which is mainly devoted to the second (hardening) phase of the evolution of SMBHBs, we will limit ourselves to an extremely simple (toy, as we call it) model, which is given solely in order to outline the contours of future detailed research. For a more detailed description of the evolution of the binary system at this phase, see, e.g. D’Orazio & Charisi (2023). This toy model consists of a set of simplifying assumptions which enable us to obtain rough but helpful understanding of properties and GW-driven evolution of potential SMBHBs. As a numerical input into our toy model and its analysis, we use just one observed value of an astrometrically oscillating source, the period of the oscillation. This single value provides a link between the observational results described in previous sections of this paper and the analysis of astrometrically oscillating sources described below. Moreover, we elaborate the toy model using one source described in this work, J21026015, as a straightforward example. The approach and corresponding physical relations described below can be applied directly to any other SMBHB system.
4.1 Formation and evolution of SMBHBs.

### v2

@@ 2 Astrometric oscillations of the quasar J02041514
The source was observed intensively by the geodetic IVS network for about 15 years until 2009. An apparently anomalous behaviour of J02041514 was reported in the ICRF2 paper (Fey et al. 2015), in which this source was referred to as a specially handled target, i.e. one requiring a special procedure for obtaining astrometric solutions. This explains its exclusion from the regular IVS observational program after 2009. The available body of VLBI observations of this source enabled us to suspect that its astrometric parameters exhibit non-random behaviour.
Bottom plots: Temporal variations of RA (left) and DEC (right) of VLBI astrometric measurements centred to the J2000.0 position RA = and DEC = (Xu et al. 2019). The trend with coefficients from Table 2 has been removed. The observational data points are shown in black with their formal error bars. An apparent decrease in data time density after 2009 is due to the recognised anomalous astrometric behaviour and exclusion of the source from regular IVS observing runs (see Section 2 for explanation and reference). The blue curves represent the model fit with parameters listed in Table 2. The shaded blue area around the blue curves indicates the 99.73% confidence interval () for the model prediction.
@@ 2.1 Brief review of astrophysical properties of the source J02041514
The object J02041514 (PKS 0202149, NRAO 91, 4C15.05) is a well-studied flat-spectrum variable radio source associated with an AGN (see a concise review of its properties by Jones et al. (2018)). Its classification is uncertain. In some works, it is referred to as a blazar (Piner & Kingham 1998), while others classify it not too reliably as a Seyfert galaxy (Véron-Cetty & Véron 2006). Its redshift has been recently estimated as (Jones et al. 2018). In the context of the current study, it is important to note that in many VLBI observations at centimetre wavelengths, the source demonstrates a core–jet morphology with apparent relative superluminal motion of its components with a speed of up to (Lister et al. 2013), where denotes the speed of light. A hint of possible transverse motion of a jet component relative to the jet axis is reported on the basis of several VLBI observations by Piner & Kingham (1998). The highest-resolution image of the source at 5 GHz obtained with the space VLBI VSOP mission confirms the jet at sub-milliarcsecond scales (Dodson et al. 2008). The source has an appreciable flux density at millimetre wavelengths (e.g. Owen et al. 1981; Chen & Wright 2009; Agudo et al. 2014). This, in combination with the astrometric oscillations discussed in the following subsection, makes the object J02041514 a suitable target for future multi-messenger studies discussed in Section 5.
@@ 2.2.1 Generalised Lomb-Scargle periodogram
| Term | Estimate | Phys. dimension | value | Estimate | Phys dimension | value |
@@
To quantitatively analyse the time series of VLBI astrometric measurements of the source J02041415, we first computed a generalised Lomb–Scargle (GLS) periodogram. The GLS method is widely employed for detecting periodicity in unevenly spaced data (Lomb 1976; Scargle 1982); see also VanderPlas (2018) for a comprehensive discussion of its practical applications in the analysis of experimental data. The technique involves fitting a periodic function, , using the (weighted) least-squares method and determining the optimal frequency, , from a predefined frequency grid that minimises the residuals. In our analysis, we adopted the periodogram formalism developed by Baluev (2014).
Let us consider a time series of measurements, , taken at times, , each associated with an RMS uncertainty, . Each observation is thus represented as a triplet (, , ), for which the periodicity (if any) is unknown. We consider two competing models for the data: the null (trend-only) model, denoted by , and the alternative model, given by
@@
and is calculated for each trial frequency, , by fitting the model to the data using the weighted least-squares method.
Several software packages are publicly available for computing the GLS periodogram; however, most of them fit either a purely periodic function or a periodic function with a constant term to the data, without accounting for linear or higher-order trends. As a result, it is often necessary to remove the trend from the data prior to computing the periodogram. Although this approach is not inherently incorrect, it may introduce a slight shift in the estimated period. Therefore, it is generally preferable to fit the periodic component simultaneously with the trend. We computed the periodogram using the GLS algorithm implemented in the Astropy library (Astropy Collaboration et al. 2013, 2018). Specifically, we used the astropy.timeseries.LombScargle class.
In right ascension (RA), the quadratic trend was found to be significant, whereas the cubic term was not. In contrast, for declination (DEC), the cubic trend was significant. Therefore, we removed the corresponding trends before computing the periodogram: the quadratic trend
@@
before computing the periodogram. By default, the fitting model in astropy.timeseries is defined as
@@ 2.2.2 Statistical significance of the astrometric oscillations
The statistical significance of the highest peak in the PSD can be expressed in terms of the false alarm probability (FAP), which quantifies the likelihood of mistaking random noise for a genuine signal. One of the most reliable methods of estimating FAP employs the bootstrap approach (Efron 1979). It evaluates the robustness of model fitting by repeatedly resampling the observed data with replacement data and refitting the model to each resample. This procedure enables the assessment of the parameter uncertainty (FAP in our case) without relying on strong distributional assumptions. As is shown by VanderPlas (2018), this approach might be more efficient than the Monte Carlo simulation in the context of PSD evaluation in the absence of reliable assumptions on the noise distribution. We generated a large number of synthetic datasets by randomly resampling the observed values with replacement data, while keeping the observation times fixed. For each resampled dataset, a GLS periodogram was computed, and the maximum peak recorded. The FAP was then estimated as the fraction of bootstrap iterations in which the maximum peak exceeds the peak observed in the original data.
Another approach involves the analytical approximation for the extreme value of FAP, developed by Baluev (2008) using the theory of stochastic processes. This method is computationally efficient and particularly advantageous when analysing large numbers of time series, such as the ones encountered in large-scale astronomical surveys. It enables the fast and reliable estimation of the FAP associated with periodogram peaks, especially when spectral leakage is minimal or can be neglected. Even in the presence of significant aliasing, the approximation remains conservative and does not increase the likelihood of false detections.
@@
| Declination: | FAP, | FAP . |
These results suggest that the detected periodicity is statistically significant in DEC, but less so in RA.
Figure 1 (upper panels, left and right) shows the GLS periodogram of the VLBI astrometry RA and DEC of J02041415, respectively. The highest significant peak in both RA and DEC occurs at the same period of years. The lower panels of Fig. 1 show the periodic fit in the time series of VLBI astrometry of J02041415. Table 2 summarises the best-fit amplitudes of trend components and harmonics together with their statistical confidence parameters, obtained in the analysis described above in subsection 2.2.1.
@@ 2.2.3 Validation of the periodicity by cross-spectrum
In order to validate the astrometric oscillation of the source J02041415 described above, we applied to the observational data the cross-spectrum method. It is used to identify periodic signals common to two time series by computing the Fourier transform of their cross-covariance function. The modulus of the cross-spectrum indicates the strength of shared periodic components; in our case, RA and DEC time series. This technique is particularly effective for detecting coherent periodicities that may be obscured in individual power spectra due to noise or aliasing (Priestley 1981).
The autocorrelation function is related to the spectral density function through the Fourier transform. Specifically, the spectral density function, , is given by
@@
where is the cross-correlation function of two time series, is the corresponding cross-spectral power density, and is the lag.
The cross-spectral density function, , is generally a complex-valued function and can be represented in two equivalent forms:
where
and are the real and imaginary parts (or symmetric and antisymmetric components) of the complex function, ,
is the amplitude spectrum, and
@@
Figure 2 shows the cross-spectrum, revealing a common period in RA and DEC of years. Although this value differs slightly from the result obtained with the GLS method in subsection 2.2.1, the difference appears to be small and could be a subject of a future detailed investigation.
@@ 3 Qualitative evaluation of the astrometric pattern of J02041514
In the case of sources with complex milliarcsecond-scale structure, absolute astrometric positions could be affected by changes in the brightness distribution with time. In extreme cases, the effect of relative fading and brightening of components could cause an apparent ‘jump’ up to mas in the radio source position (Titov et al. 2022), (Osetrova et al. 2024). However, generating a persistent harmonic oscillating celestial trajectory this way is very unlikely.
@@
The oscillating pattern of absolute celestial co-ordinate estimates for the source J02041514 might have a variety of explanations. We must note that VLBI astrometry deals with centroids of the source’s brightness distribution, which are usually associated with the base of the jet in AGNs. These are optically thick regions of jets with typically flatter or inverted spectra at radio frequencies. The jet base areas do not coincide with the 3D position of central SMBHs. Thus, what is seen in Fig. 1 is not the motion of the SMBHBs themselves, but their motion manifesting itself in the apparent motion of the astrometrically detected structural component of the source. A detailed physical description of the causal relationship between the binary system’s orbital period and the observed period of astrometric oscillations associated with a jet-related morphological component warrants a dedicated investigation, which lies beyond the scope of the present paper. However, recent general relativistic magnetohydrodynamic (GRMHD) simulations of binary black hole systems suggest that, depending on the system’s eccentricity, the periodicity of accretion and jet emission ranges from approximately 1.0 to 0.7 times the binary orbital period (Manikantan et al. 2025). Therefore, in the subsequent sections, our assumption that the period of the SMBHB in J02041514 is equal to the observed period of astrometric oscillations introduces no significant error for the toy model considered.
@@
Various effects might be invoked for explanations of the apparent oscillating behaviour of the jet base. One of the most studied cases, involving both long-term flux density variability monitoring in optics and radio, as well as structural studies at VLBI (milliarcsecond) scales, is the blazar OJ287, considered to be a likely SMBHB. However, in this and several other similar cases, alternative physical causes of temporal, structural, or astrometric manifestations, such as the impact of Lense–Thirring effect on an accretion disc or jet precession, are considered to be responsible for the observed phenomenology (for a review of the issue see Britzen et al. (2023) and references therein). Among those causes, periodic quasi-orbital motion of the jet base associated with an orbital motion of SMBH in a binary system is at least no less plausible than others, especially coupled with the modality of SMBHB existence in the Universe (see Section 1). Moreover, SMBHBs must evolve, and the apparent orbital motion of an AGN, detected via oscillating behaviour of associated with bright morphological features in VLBI images of AGNs, could be looked at as an indicator of a potential presence of a SMBHB.
@@
To a certain extent, our interpretation of an apparent oscillating astrometric pattern detected in the radio domain is similar to varstrometric detections of potential sub-kiloparsec dual AGNs based on Gaia optical astrometry (Hwang et al. 2020). Another search for potential SMBHBs via their quasi-periodic patterns in optical light curves was reported recently by Chen & Jiang (2024). We also note the recent work by Makarov et al. (2024) in which a novel periodogram analysis algorithm applied to VLBI astrometric monitoring of 259 ICRF3 sources resulted in detection of 49 objects with quasi-periodic celestial position variations, including the source in which we are interested, J02041514. They report the detection of a quasi-periodic signal in DEC only, with a period of 8.05 years. This value slightly differs from our period estimate of 7.9 years. The reason for this discrepancy requires further investigation. We note, however, that, qualitatively, the absence of a detected quasi-periodic signal in RA for J02041514 by Makarov et al. (2024) is consistent with the clearly lower confidence in periodicity detection in RA compared to DEC in our analysis (see Fig. 1 and FAP values in Table 2 and subsection 2.2.2).
@@
Additional justification of our logic comes from the fact that if an SMBHB evolves all the way towards full coalescence, accompanied by a GW burst, it must end up as a single SMBH. Indeed, some known objects such as the quasar 3C 186 are suspected to be post-coalescence products of a former SMBHB (Boschini et al. 2024). Another suspected case of merging SMBHs or even a post-coalescence recoiling SMBH is reported recently in the radio quasar PKS 090357 based on optical spectroscopy (Goldoni et al. 2024).
@@
In the following section, we analyse the evolution of SMBHBs with astrometric oscillations. Such a binary system will eventually become a multi-messenger with very distinct electromagnetic and GW imprints. We do so for the source discussed in this work, J02041514. This source offers a straightforward example, allowing us to focus on properties of SMBHB as a GW generator. However, the analysis presented in the following section is applicable to other sources with apparent astrometric oscillations, and wider SMBHBs identified by other techniques.
@@ 4 Toy model of SMBHB evolution
As was mentioned in the introduction, the formation and evolution of SMBHBs encompasses three phases: the pairing phase, the hardening stage, and the final coalescence phase (see Begelman et al. 1980; Colpi 2014, and references therein). In this section, which is mainly devoted to the second (hardening) phase of the evolution of SMBHBs, we limit ourselves to an extremely simple (toy, as we call it) model, which is given solely in order to outline the contours of future detailed research (for a more detailed description of the evolution of the binary system at this phase, see e.g. D’Orazio et al. (2024)). This toy model consists of a set of simplifying assumptions that enable us to obtain a rough but helpful understanding of properties and GW-driven evolution of potential SMBHBs. As a numerical input into our toy model and its analysis, we use just one observed value of an astrometrically oscillating source, the period of the oscillation. This single value provides a link between the observational results described in previous sections of this paper and the analysis of astrometrically oscillating sources described below. Moreover, we elaborate the toy model using one source described in this work, J21026015, as a straightforward example. The approach and corresponding physical relations described below can be applied directly to any other SMBHB system.
4.1 Formation and evolution of SMBHBs